OpenStudy (anonymous):

Given a right triangle, one leg is 10 in. and the hypotenuse is 15 in., find the length of the other leg using the Pythagorean theorem. Write your answer as a square root, not a decimal. (Use sq rt for the square root symbol).

OpenStudy (anonymous):

We know that the Pythagorean Theorem is:\[c^{2}=a^{2}+b^{2}\] Since we know 'a' and 'c' we need to solve for 'b':\[b=\sqrt{c^{2}-a^{2}}\] Now we can fill in know values and solve.

OpenStudy (anonymous):

I dont know how to do the rest could you help me !!:)

OpenStudy (anonymous):

Sure. We know that 'c' is 15 and 'a' is 10 so: \[b=\sqrt{15^{2}-10^{2}}\] Can you go from here?

OpenStudy (anonymous):

okay so 15(2) is 225 and 10(2) is 100 is the correct ?

OpenStudy (anonymous):

Yes! Good so far

OpenStudy (anonymous):

okay whats next ? do I subtract them ? and if yes how ?

OpenStudy (anonymous):

So now you have:\[b=\sqrt{225-100}\]. If you subtract you get \[b=\sqrt{125}\]

OpenStudy (anonymous):

This is usually good enough for the answer or you can simplify if required.

OpenStudy (anonymous):

okay so i sqaure rooted 125 and i got 11 so is that goung to be my answer ?

OpenStudy (anonymous):

\[\sqrt{125}=5\sqrt{5} \approx 11.18\] All of them are correct, but the original question askes you to leave it in root form. So, the first one.

OpenStudy (anonymous):

so 125

OpenStudy (anonymous):

No not just 125. Square root 125 \[\sqrt{125}\]

OpenStudy (anonymous):

ok so like this \[\sqrt{125}

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

You're welcome.